OpenStudy (anonymous):

solve for initial problem of: dy/dx: x(2+x^2)^4, y(0)=0

OpenStudy (anonymous):

just integrate the function. the constant of integration can be taken as y(0)=0.

OpenStudy (anonymous):

I did try to ingrate the function but I think I did it wrong or something

OpenStudy (anonymous):

take u=2+x^2 then, du=2xdx so,\[\frac{1}{2}\int\limits_{}^{}u^4 du\]go from there

OpenStudy (anonymous):

But, then y(0) would not equal zero...so perhaps I am misunderstanding the question?

OpenStudy (anonymous):

I got it, thanks

OpenStudy (anonymous):

Unless it should read\[y_0=0\]?

OpenStudy (anonymous):

can you help me on something else?

OpenStudy (anonymous):

no, it's y(0)=0

OpenStudy (anonymous):

that makes no sense. y(0)=32/10 not 0

OpenStudy (anonymous):

idk, the solution is y=1/10(2+x^2)^5-16/5

OpenStudy (anonymous):

so I guess the answer is:\[y(x)=\frac{1}{10}(2+x^2)^5-\frac{32}{10}\]

OpenStudy (anonymous):

lol...yeah same thing.

OpenStudy (anonymous):

how did you get that?? that is right

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

can you walk me thru that?

OpenStudy (anonymous):

what did you get for the integral?

OpenStudy (anonymous):

should I multiply out first before integral?

OpenStudy (anonymous):

you get\[u=\frac{1}{10}u^5+u_0\]for the integral

OpenStudy (anonymous):

then sub back in u=2+x^2 \[y(x)=\frac{1}{10}(2+x^2)^5+y_0\]Since we know y(0)=0...we can solve for y_0

OpenStudy (anonymous):

oh ic, thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can you help me on something else?

OpenStudy (anonymous):

ok. post it as a new question tho

OpenStudy (anonymous):

thanks bye in here